OpenStudy (anonymous):

Let f:Rn -> Rn and g:Rn -> Rn be inverse functions. Suppose f and g are differentiable. Show that determianants of Df and Dg are not zero, and that Dg=(Df)^-1. Where Df and Dg is the derivative of f and g respectively.

OpenStudy (cruffo):

The "Df and Dg are not zero" part should come directly from the definition of differentiable. However, I don't see how "Dg=(Df)^-1" could be possible in general. Are you given any more info on f and g?

OpenStudy (cruffo):

Ahhh.. f and g are inverses of each other...

OpenStudy (anonymous):

(Df)^-1 is the inverse matrix of Df btw

OpenStudy (cruffo):

yep. I know the notation.

OpenStudy (anonymous):

that is all the info given and I'm not sure how the not zero part follows

OpenStudy (anonymous):

i mean i understand that Df and D(f^-1)=Dg exist but unless I can show D(f^-1) = (Df)^-1 then i cannot say that since they are invertible the determinant is nonzero...

OpenStudy (cruffo):

First paragraph of the following http://people.math.gatech.edu/~ghomi/LectureNotes/LectureNotes6G.pdf may provide the idea of how to prove that the Jacobian is nonsingular. I'll look back at my DG book.

OpenStudy (anonymous):

thx for the tip, gona take a look now....

OpenStudy (cruffo):

I got the following Since \(f \circ g = g \circ f = I(R_n)\), then \(D(f \circ g) = Df_{|g} \cdot Dg = D I(R_n)\). So \(|Df_{|g} \cdot Dg | = |Df_{|g}| \cdot |Dg | = |D I(R_n)|\). But \( |D I(R_n)| = 1\). Thus \(|Dg | \neq 0\). Similiar proof for \(|Df|\).